Arrange the following growth functions in increasing order of growth:

       O(n1.5), O(2n), O (nlogn), O(logn), O(n3),  O() ,O(),O()


1
Expert's answer
2021-11-03T17:31:08-0400

The given growth function,

"O(n^{1.5}), O(2n), O (nlogn), O(logn), O(n^3)"

Again,

"O(n^{\frac{3}{2}}), O(2n), O (nlogn), O(logn), O(n^3)"

Now,

As we know that, big O notation provides the upper bounds to the function,

"O(n^3)>O(n^{\frac{3}{2}})>O(2n)>O(\log n^n)>O(\log n)"